OpenStudy (anonymous):

Neutron stars are extremely dense objects that are formed from the remnants of supernova explosions. Many rotate very rapidly. Suppose the mass of a certain spherical neutron star is twice the mass of the Sun and its radius is 5.00 km. Determine the greatest possible angular speed the neutron star can have so that the matter at its surface on the equator is just held in orbit by the gravitational force.

OpenStudy (anonymous):

mass of sun is 1.98892 x 10^{30} kg mass of neutron star = 2x1.98892 x 10^{30} kg=3.97784 x 10^{30} \[g = \frac{(6.67300 × 10^{-11})(3.97784 \times 10^{30})}{(5000)^2}=\frac{v^2}{5000}=a\] Now, solve for v.

OpenStudy (anonymous):

why 5000?

OpenStudy (anonymous):

that's the radius of the star